OpenStudy (anonymous):

OH MY GOSH!!!! SOMEONE PLEASE HELP, C'MON ISN'T THIS WHAT THE WEBSITE IS FOR ????????????????? ]:(

OpenStudy (anonymous):

A picture frame is shaped as a heptagon. The measure of one angle is 216°. The remaining interior angles are congruent. What is the measure of each remaining interior angle in the picture frame? 186° 150° 114° 180°

OpenStudy (anonymous):

no i'm in geometry at florida virtual school and it lesson 6

OpenStudy (anonymous):

it's sad how i have to do all this to get a lil attention.. smh

Parth (parthkohli):

The number of sides and angles in a heptagon is \(7\). The sum of all interior angles is \(180(n - 2)\) where \(n = 7\).

OpenStudy (anonymous):

so the answer is 180 you say?

Parth (parthkohli):

Determine \(180 \times (7 - 2)\). Now the sum of all angles will be \(216 + n + n + n + n + n + n = 180\times(7 - 2)\) where \(n\) is the angle you are looking for.

OpenStudy (anonymous):

ok.so 180 then ?

Parth (parthkohli):

Did you solve the equation for \(n\)?

OpenStudy (anonymous):

huh? no i don't get it.

Parth (parthkohli):

\[216 + 6n = 180\times(7-2)\]Solve the equation.

OpenStudy (hba):

Solve for n ^

OpenStudy (anonymous):

i got n equal 6, that doesnt even look right

Parth (parthkohli):

Do you think \(216 + 6\cdot6 = 180\times(7-2)\)

OpenStudy (anonymous):

ok so i kinda get what your saying with the 216+6.6 = 180 but what about x (7-2) oh my gosh i'm so confused :/

Parth (parthkohli):

What is \(7 - 2\)?

Parth (parthkohli):

\(\times\) means "times".

OpenStudy (anonymous):

5 and i know that but how do you work all that out with the rest

Parth (parthkohli):

What is \(180\) times \(5\)?

OpenStudy (anonymous):

900